OpenStudy (anonymous):

log8 (x+1) = log8 (2x-2) log2 x+log2 (x+4) = 5 log4 (x-2) = 2

OpenStudy (anonymous):

e^(-3x+4) =22 ln(e^(-3x+4)) = ln(22) // ln(e^y) = y -3x+4 = ln(22) -3x = ln(22)-4 x = (ln(22)-4)/-3

OpenStudy (anonymous):

how many problems is this?

OpenStudy (anonymous):

four

OpenStudy (anonymous):

....now three.

OpenStudy (anonymous):

\[\log_4 (x-2) = 2\iff 4^2=x-2\] solve for \(x\)